# The functional Loomis-Whitney type inequality in the Heisenberg groups and Projection theorems over finite fields

### v1

@@ 2.3 Counting subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
$\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\leq\prod_{j=1}^{2n-1}\|f_{j}\|_{L^{\ell_{j}}(\mathbb{F}^{2n}_{q},dx)}$ (30)
@@
To prove (30), we apply the discrete version of the multilinear interpolation theorem, a generalization to [17, Proposition 6.1]. Since $(\mathbb{F}_{q}^{2n},dx)$ is a finite measure space equipped with the normalized counting measure, the same interpolation argument carries over verbatim. We then perform (n1)(n-1) iterations with $\theta=\tfrac{1}{k}$, starting from $k=2$ and ending at $k=n$, using the identity
@@
where each $K_{j}$ is viewed as a subset of $\mathbb{H}^{n}(\mathbb{F}_{q})$ through the identification $\mathbb{F}_{q}^{2n}=W_{j}\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j=1,\ldots 2n$. Then for $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q}),$ we can write
@@
Here all $L^{p}$-norms are taken with respect to the normalized counting measure on $\mathbb{F}_{q}^{2n}$, so that
$\|\mathbf{1}_{K_{j}}\|_{L^{p}}=|K_{j}|^{1/p},\qquad j=1,\dots,2n.$
Finally, applying Theorem 1.4 with $f_{j}=\chi_{K_{j}}$ for $j=1,\ldots,2n$ immediately yields (34), as desired. ∎
@@
To prove Theorem 1.6, we invoke the following incidence bound, due to Vinh [23].
@@
Let $\mathscr{P}$ be a set of points in $\mathbb{F}_{q}^{2}$ and $\mathscr{L}$ a set of lines in $\mathbb{F}_{q}^{2}$. The number of incidences between $\mathscr{P}$ and $\mathscr{L}$, denoted by $I(\mathscr{P},\mathscr{L})$, satisfies
@@
Let $f_{1}$ and $f_{2}$ be the indicator functions of $\pi_{1}(K)$ and $\pi_{2}(K)$, respectively. Note that the equation $a_{1}b_{1}+b_{2}=a_{2}$ is equivalent to an incidence between a point $(a_{1},a_{2})$ and a line defined by $y=b_{1}x+b_{2}$. Let $\mathscr{P}$ and $\mathscr{L}$ be the corresponding sets of points and lines, i.e. $|\mathscr{P}|=|\pi_{1}(K)|$ and $|\mathscr{L}|=|\pi_{2}(K)|$. Therefore, applying Theorem 6.1, one has
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{F}_{q}^{3}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{|K|}{q^{3}}\leq\frac{I(\pi_{1}(K),\pi_{2}(K))}{q^{3}},$
which is at most
$\frac{|\pi_{1}(K)||\pi_{2}(K)|}{q^{4}}+2q^{-\frac{5}{2}}\sqrt{|\pi_{1}(K)||\pi_{2}(K)|}.$
This implies the desired conclusion. ∎
@@ 7 Acknowledgements
Thang Pham and Dung The Tran would like to thank to the VIASM for the hospitality and for the excellent working condition. Daewoong Cheong sincerely thanks the Korea Institute for Advanced Study for financial support and hospitality and was partially supported by the research year program of Chungbuk National University in 2025.

### v2

@@ 2.3 Counting subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
$\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\lesssim\prod_{j=1}^{2n-1}\|f_{j}\|_{L^{\ell_{j}}(\mathbb{F}^{2n}_{q},dx)}$ (26)
@@
To prove (26), we apply the discrete version of the multilinear interpolation theorem, a generalization to [14, Proposition 6.1]. Since $(\mathbb{F}_{q}^{2n},dx)$ is a finite measure space equipped with the normalized counting measure, the same interpolation argument carries over verbatim. We then perform (n1)(n-1) iterations with $\theta=\tfrac{1}{k}$, starting from $k=2$ and ending at $k=n$, using the identity
@@
where each $K_{j}$ is viewed as a subset of $\mathbb{H}^{n}(\mathbb{F}_{q})$ through the identification $\mathbb{F}_{q}^{2n}=W_{j}\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j\in\{1,\ldots 2n\}$. Then for $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q}),$ we can write
@@
Note that
$\|\mathbf{1}_{K_{j}}\|_{L^{p}}=\bigg(\frac{|K_{j}|}{q^{2n}}\bigg)^{1/p},\qquad j=1,\dots,2n.$
Finally, applying Theorem 1.4 with $f_{j}=\mathbf{1}_{K_{j}}$ for $j\in\{1,\ldots,2n\}$, multiply both sides by $q^{2n+1}$ and simplify the power of $q$ to obtain (30), as desired. ∎
@@
Let $\mathscr{P}$ be a set of points in $\mathbb{F}_{q}^{2}$ and $\mathscr{L}$ be a set of lines in $\mathbb{F}_{q}^{2}$. The number of incidences between $\mathscr{P}$ and $\mathscr{L}$, denoted by $I(\mathscr{P},\mathscr{L})$, is defined by
$I(\mathscr{P},\mathscr{L})=|\{(p,\ell)\in\mathscr{P}\times\mathscr{L}\colon p\in\ell\}|.$
To prove Theorem 1.6, we invoke the following incidence bound due to Vinh [20].
@@
Let $\mathscr{P}$ be a set of points in $\mathbb{F}_{q}^{2}$ and $\mathscr{L}$ a set of lines in $\mathbb{F}_{q}^{2}$. The number of incidences between $\mathscr{P}$ and $\mathscr{L}$ satisfies
@@
Let $f_{1}$ and $f_{2}$ be the indicator functions of $\pi_{1}(K)$ and $\pi_{2}(K)$, respectively. Each element $(x_{1},x_{2},t)\in K$ determines a point $(a_{1},a_{2})=\pi_{1}(x_{1},x_{2},t)\in P$ and a line $\ell=\pi_{2}(x_{1},x_{2},t)\in L$, which satisfy the relation $a_{1}b_{1}+b_{2}=a_{2}$. Hence each element of $K$ gives rise to an incidence between $P$ and $L$.
Moreover, when $n=1$, the map
$(x_{1},x_{2},t)\longmapsto\bigl(\pi_{1}(x_{1},x_{2},t),\,\pi_{2}(x_{1},x_{2},t)\bigr)$
is injective, so distinct elements of $K$ yield distinct incidences. Therefore,
$|K|\leq I(P,L).$
Applying Theorem 6.1, we obtain
$\frac{|K|}{q^{3}}\leq\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q^{4}}+2q^{-5/2}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}.$
This yields the desired conclusion. ∎
7 Connections to other topics
In this section, we address how our results are connected to some questions in the literature.
7.1 Bhowmik, Iosevich, Koh, and Pham’s boundedness problem
Let $\mathcal{K}\colon\mathbb{F}_{q}^{d}\times\mathbb{F}_{q}^{d}\to\mathbb{R}$ be any function (which, below, will play the role of the kernel of an integral operator for the counting measure), and $\mathcal{G}=(V,E)$ a connected ordered graph on $m$ vertices. In a recent paper [4], Bhowmik, Iosevich, Koh, and Pham introduced the boundedness problem for a multilinear operator associated with $\mathcal{K}$ and $\mathcal{G}$. To be precise, define
by
The operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ is a generalization of several geometric counting functions:
  1. 1.
  2. 2.
  3. 3.
    If $\mathcal{G}$ is a tree and $f_{1}=\cdots=f_{m}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of copies of the tree $\mathcal{G}$ in $A$.
The inequality (7) for $n=1$ is essentially the same as the inequality (31) for $m=2$, with suitable $\mathcal{G}$ and $\mathcal{K}$ chosen. To see this, define $\phi\colon\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ by $\phi(x,y,t)=(x,y,t+\frac{xy}{2})$. Then, since $\phi$ is a bijective map, the left-hand side of (7) can be written
Then the left-hand side $\Lambda_{\mathcal{G}}^{\mathcal{K}}(f_{1},f_{2})$ of the inequality (31) is equal to
which coincides with the value (7.1). Thus, in this case, both inequalities (7) and (31) coincide with each other. This connection suggests a further study of the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ for various $\mathcal{K}$ and $\mathcal{G}$.
7.2 The orthogonal projection problem
For a subset $K\subset\mathbb{F}_{q}^{2n+1}$ and an integer $r\geq 1$, let
$\mathscr{E}_{r}=\mathscr{E}_{r}(K)\;:=$ $\;\big\{\,W\leq\mathbb{F}_{q}^{2n+1}:\ \dim W=2n,\;\;K\text{ can be covered by at most }r$
 $~~~~\text{ (additive) translates of }W^{\perp}\,\big\},$
where a translate of $W^{\perp}$ means a subset of $\mathbb{F}_{q}^{2n+1}$ of the form $(x,t)+W^{\perp}$ for some $(u,\tau)\in\mathbb{F}_{q}^{2n+1}$ Chen [6] proved that
Informally, this states that for most subspaces $W\subset\mathbb{F}_{q}^{2n+1}$, the covering number of $K$ by translates of $W^{\perp}$ is not very small.
Chen’s result can be interpreted from the view of the (orthogonal) projection problem. To see this, for each $j=1,\ldots,2n$, let $P_{j}:\mathbb{F}_{q}^{2n+1}\mapsto W_{j}$ be the orthogonal projection defined by $P_{j}(x,t):=(\check{x}_{j},t),$ where $W_{j}$ is the coordinate hyperplane given in (3). Note that $W_{j}^{\perp}=\mathrm{Ker}\,P_{j}$ and the translates of $W_{j}^{\perp}$ are exactly cosets of the subspace $W_{j}^{\perp}$ in $\mathbb{F}_{q}^{2n+1},$ which are in turn the fibres of $P_{j}$. Note that each translate of $W_{j}^{\perp}$ can be written in the form $(u,\tau)+W_{j}^{\perp}$ for uniquely determined $(u,\tau)\in W_{j}.$ Thus, we observe that for a subset $K$ of $\mathbb{F}_{q}^{2n+1}$, $K\cap((u,\tau)+W_{j}^{\perp})\neq\emptyset$ for $(u,\tau)\in W_{j}$ if and only if $(u,\tau)\in P_{j}(K)$. Thus, one can write
$K\;\subset\;\bigcup_{(u,\tau)\in P_{j}(K)}\left((u,\tau)+W_{j}^{\perp}\right).$ (34)
This implies that the minimal number of translates of $W_{j}^{\perp}$ that cover a set $K\subset\mathbb{F}_{q}^{2n+1}$ coincides with $|P_{j}(K)|.$
Now let $\mathscr{H}=\{W_{1},\dots,W_{2n}\}$ be the set of coordinate hyperplanes in $\mathbb{F}_{q}^{2n+1}$. If $|\mathscr{H}|=2n>|\mathscr{E}_{r}|$, then, by the definition of $\mathscr{E}_{r}$, there exists at least one coordinate hyperplane $W_{j}\in\mathscr{H}$ that does not belong to $\mathscr{E}_{r}$, and for this $j$ the covering number of $K$ by translates of $W_{j}^{\perp}$ is strictly larger than $r$, i.e., $P_{j}(K)$ contains more than $r$ points.
Combining this observation with Chen’s estimate (33) and the fact that $|\mathscr{H}|=2n$, one finds a constant $c_{n}>0$ depending only on $n$ such that for every nonempty set $K\subset\mathbb{F}_{q}^{2n+1}$ there exists an index $j_{0}\in\{1,\dots,2n\}$ with
$|P_{j_{0}}(K)|\;\geq\;c_{n}\,\frac{|K|}{q^{2n}}.$ (35)
Thus, among the given $2n$ projections $P_{j}:\mathbb{F}_{q}^{2n+1}\mapsto W_{j}$ for $j=1,\ldots,2n,$ there exists at least one projection $P_{j_{0}}$ under which the image of $K$ must have the size comparable to $|K|/q^{2n}$.
This discussion can be phrased in a more general framework. More precisely, let $(H,\cdot)$ be a (finite) set equipped with an algebraic structure with the identity and let $\{\Phi_{i}:H\mapsto V_{i}\,|\,i\in I\}$ be a family of algebraic projections, where $V_{i}$ are subobjects of $H$ with the same cardinality. Let $Z_{i}:=Ker\,\Phi_{i}$. Then the situation is analogous to the case of the orthogonal projections $P_{j}$: Each coset of $Z_{j}$ can be expressed in the form $v\cdot Z_{j}$ for uniquely determined $v\in V_{j},$ and a coset $v\cdot Z_{j}$ for $v\in V_{j}$ is equal to the fiber $\Phi_{i}^{-1}(v).$ Also we have that for a subset $K$ of $H$, $K\cap v\cdot Z_{i}\neq\emptyset$ for $v\in V_{i}$ if and only if $v\in\Phi_{i}(K)$. Thus, one can write
$K\;\subset\;\bigcup_{v\in\Phi_{i}(K)}v\cdot Z.$ (36)
In particular, the minimal number of cosets of $Z_{i}$ in $H$ that cover $K$ is equal to $|\Phi_{i}(K)|.$ In the light of (35), one seek an index $i_{0}\in I$ for which the image $\Phi_{i_{0}}(K)\subset V_{i_{0}}$ is as large as possible. In practice, depending on which features of the original set one wishes to preserve, the choice of an appropriate family of projection maps is crucial. Our case concerns $H=\mathbb{H}^{n}(\mathbb{F}_{q})$ and $\Phi_{i}=\pi_{i}$ for $i\in I=\{1,\ldots,2n\}.$
In the following, we compare Chen’s theorem and Corollary 1.5 from this point of view. For each $j=1,\ldots,2n$, define $T_{j}:\mathbb{H}^{n}(\mathbb{F}_{q})\to\mathbb{F}_{q}^{2n+1}$ by
This map is a bijective map with $P_{j}\circ T_{j}=\pi_{i}$ for all $j=1,\ldots,2n$. We observe that for each $(u,\tau)\in W_{j}$, we have
$T_{j}\big(\pi_{j}^{-1}(u,\tau)\big)=P_{j}^{-1}(u,\tau),$ (38)
i.e., $T_{j}$ maps each Heisenberg fibre $\pi^{-1}(u,\tau)$ bijectively onto the fibre $P_{j}^{-1}(u,\tau)$. Moreover, the restriction of $T_{j}$ to $W_{j}$ is the identity map.
Let $K\subset\mathbb{\mathbb{}}H^{n}(\mathbb{F}_{q})$ and fix $j=1,\ldots,2n.$ Recall that $\mathrm{Ker}\,\pi_{j}=W_{j}^{\perp}$. Letting $H=\mathbb{H}^{n}(\mathbb{F}_{q})$ and $\Phi_{j}=\pi_{j}$, by (36) we have
$K\;\subset\;\bigcup_{(u,\tau)\in\pi_{j}(K)}(u,\tau)\cdot W_{j}^{\perp}.$ (39)
Applying $T_{j}$ on both sides of (39), by the equality (38) and the facts that $\pi_{j}^{-1}(u,\tau)=(u,\tau)\cdot W_{j}^{\perp}$ and $P_{j}^{-1}(u,\tau)=(u,\tau)+W_{j}^{-1}$, we obtain
$T_{j}(K)\;\subset\;\bigcup_{(u,\tau)\in\pi_{j}(K)}\bigl((u,\tau)+W_{j}^{\perp}\bigr).$
Note that each coset $(u,\tau)+W_{j}^{\perp}$ in the above union is necessary to cover $T_{j}(K).$ Thus, the quantity $|\pi_{j}(K)|$ coincides with the covering number of $T_{j}(K)$ by translates of $W_{j}^{\perp}$. Now for each $j=1,\ldots,2n$, define
$\phi_{j}\;:=\;T_{j}^{-1}\circ P_{j}\circ T_{j}:\mathbb{H}^{n}(\mathbb{F}_{q})(\cong\mathbb{F}_{q}^{2n+1})\to\mathbb{H}^{n}(\mathbb{F}_{q})(\cong\mathbb{F}_{q}^{2n+1}).$
These maps enjoy three basic features that we use repeatedly: they are idempotent retractions,
$\phi_{j}\circ\phi_{j}=\phi_{j};$
they have the same fibres as $\pi_{j}$, in the sense that $\phi_{j}(x)=\phi_{j}(y)$ if and only if $\pi_{j}(x)=\pi_{j}(y)$; and, consequently, for every subset $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$,
$|\phi_{j}(K)|=|\pi_{j}(K)|=\bigl|P_{j}\bigl(T_{j}(K)\bigr)\bigr|.$
Corollary 1.5 yields a constant $C_{n}>0$ such that
$\max_{1\leq j\leq 2n}|\pi_{j}(K)|\;\geq\;C_{n}\,|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}},$
and hence there exists an index $j^{\prime}\in\{1,\dots,2n\}$ for which
$|\phi_{j^{\prime}}(K)|\;=\;|P_{j^{\prime}}(T_{j^{\prime}}(K))|\;\geq\;C_{n}\,|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}}.$ (40)
We now compare between (35) and (40). Using only the trivial constraint $|K|\leq q^{2n+1}$, one checks that
$\frac{|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}}}{|K|/q^{2n}}\;\geq\;q^{2n-1}$
for all nonempty $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})(\cong\mathbb{F}_{q}^{2n+1})$.
Thus, with the same $\mathscr{H}$, the family $\{\pi_{j}\}_{j}$ of projections provides a strictly stronger lower bound for the maximal projection size in one of the vertical directions.
Another related question is a discrete analogue to the problem considered by Balogh, Fässler, Mattila, and Tyson [2] for the real Heisenberg group. To explain this, for $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ and $1\leq m\leq n$, let $\mathscr{E}_{r}^{h,m}(K)$ denote the set of homogeneous subgroups $G\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ of size $q^{m}$ such that $K$ can be covered by at most $r$ left cosets of $G^{\perp}$. (See Section 2 for the notion homogeneous and the notation $G^{\perp}$.) Then following Balogh, Fässler, Mattila, and Tyson, it is natural to seek upper bounds for $|\mathscr{E}_{r}^{h,m}(K)|$ in terms of $|K|$, $q$, $n$, and $m$. We plan to pursue this Heisenberg covering problem in a sequel paper.
@@ 8 Acknowledgements
Thang Pham and Dung The Tran would like to thank VIASM for the hospitality and for the excellent working conditions. Daewoong Cheong sincerely thanks the Korea Institute for Advanced Study for financial support and hospitality and was partially supported by the research year program of Chungbuk National University in 2025.